OpenStudy (anonymous):

solve for the equation, smaller and larger values (x+1)^4+8=246

OpenStudy (anonymous):

subtract 8 take the fourth root subtract 1

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

By subtracting 8 from both sides, then taking the fourth root of both sides, and finally subtracting 1 from both sides, you are left with \[x =\left(\begin{matrix}+ \\-\end{matrix}\right)\sqrt[4]{238} -1\] These are the exact values of x, but if you're looking for an approximation (to three decimal places), x = 2.928 or x = -4.928

OpenStudy (anonymous):

^That was the perfectly correct answer, but is it acceptable to just give answers out when the process to solve the problem is outlined? New to this site, and just wondering what is generally acceptable here.